Fred is 3 years older than George. In 4 years from now, Fred will be three times as old as George was 5 years ago. How old are they now?
Answer by alicealc(293)

Fred's age = f
George's age = g
f = g + 3
4 years from now Fred will be three times as old as George was 5 years ago:
f + 4 = 3 * (g - 5)
f + 4 = 3g - 15
g + 3 + 4 = 3g - 15
g + 7 = 3g - 15
7 + 15 = 3g - g
22 = 2g
22/2 = g
11 = g
f = g + 3 = 11 + 3 = 14

so, Fred is 14 years old and George is 11 years old now